OpenStudy (anonymous):

how can you solve 13/16x + 5/4 = 2/8x + 4 ?

OpenStudy (anonymous):

Is it the fraction times x or is x in the denominator (bottom)

OpenStudy (mertsj):

\[\frac{13}{16}x+\frac{5}{4}=\frac{2}{8}x+4\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (mertsj):

The asker has left the building.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

x is the variable

OpenStudy (texaschic101):

I would start by multiplying in the equation by 16. This will get rid of the fractions

OpenStudy (anonymous):

Like mertsj wrote it

OpenStudy (texaschic101):

typo...multiplying everything in the equation by 16

OpenStudy (texaschic101):

13/16x(16) + 5/4(16) = 2/8x(16) + 4(16) 13x + 20 = 4x + 64 -- subtract 20 from both sides 13x = 4x + 64 - 20 -- subtract 4x from both sides 13x - 4x = 64 - 20 -- combine like terms 9x = 44 x = 44/9

OpenStudy (texaschic101):

any questions ?

OpenStudy (texaschic101):

if you multiply the equation by the LCD of all the terms, then it will get rid of the fractions

OpenStudy (anonymous):

ok thanks